# Temperature is All You Need for Generalization in Langevin Dynamics and other Markov Processes

### v1

@@ F.1 Experimental results
The following are results of training with SGLD (a discretized version of the CLD dynamics in (10)) on a few benchmark datasets. We use several different values of $\beta$ relative to $N$ (number of training samples). For simplicity, we focused on binary classification cases. In all datasets with more than 2 classes, we constructed a binary classification task by partitioning the original labels set into $2$ disjoint sets of the same size.
The results demonstrate that learning with SGLD is possible with various values of $\beta$. In fact, in several instances, the injected noise appears to improve the generalization gap, e.g, in SVHN, in all the tested $\beta$ values between $0.4\cdot N$ and $2\cdot N$ the average test error remained almost the same while the training error decreased as $\beta$ increased (i.e. the generalization gap increased). Notably, we also observe that for sufficiently large levels of noise the generalization bounds are non-vacuous.
@@
$0.03\cdot N$ $0.1161(\pm 0.0028)$ $0.1074(\pm 0.0035)$0.0087 0.10461
$0.15\cdot N$ $0.0497(\pm 0.0014)$ $0.0494(\pm 0.0031)$0.0003 0.23359
$0.4\cdot N$ $0.0281(\pm 0.0002)$ $0.0358(\pm 0.0029)$ 0.0077 0.38137
$0.7\cdot N$ $0.0202(\pm 0.0006)$ $0.0284(\pm 0.0024)$ 0.0082 0.50448
$N$ $0.0162(\pm 0.0006)$ $0.0278(\pm 0.0023)$ 0.0116 0.60296
$2\cdot N$ $0.0092(\pm 0.0004)$ $0.0262(\pm 0.0016)$ 0.017 0.85268
$\infty$ $0.0001(\pm 0)$ $0.0229(\pm 0.0004)$ 0.0228 $>1$
@@
$0.03\cdot N$ $0.0999(\pm 0.001)$ $0.1087(\pm 0.0167)$ 0.0088 0.11705
$0.15\cdot N$ $0.0765(\pm 0.0009)$ $0.0803(\pm 0.0015)$ 0.0038 0.26145
$0.4\cdot N$ $0.0635(\pm 0.0005)$ $0.0722(\pm 0.002)$ 0.0087 0.42686
$0.7\cdot N$ $0.0567(\pm 0.0006)$ $0.0691(\pm 0.0019)$ 0.0124 0.56466
$N$ $0.0525(\pm 0.0005)$ $0.0675(\pm 0.0013)$ 0.015 0.67489
$2\cdot N$ $0.043(\pm 0.0007)$ $0.0672(\pm 0.0023)$ 0.0242 0.95442
$\infty$ $0.0248(\pm 0.001)$ $0.0675(\pm 0.0033)$ 0.0427 $>1$
@@
$0.01\cdot N$ $0.0746(\pm 0.0012)$ $0.1033(\pm 0.0032)$ 0.0287 0.0589
$0.03\cdot N$ $0.0441(\pm 0.0004)$ $0.067(\pm 0.0026)$ 0.0229 0.10198
$0.15\cdot N$ $0.0251(\pm 0.0005)$ $0.0445(\pm 0.002)$ 0.0194 0.22801
$0.4\cdot N$ $0.0182(\pm 0.0005)$ $0.0374(\pm 0.0017)$ 0.0192 0.37233
$0.7\cdot N$ $0.0146(\pm 0.0004)$ $0.0363(\pm 0.002)$ 0.0217 0.49255
$N$ $0.0124(\pm 0.0002)$ $0.0342(\pm 0.0014)$ 0.0218 0.58871
$2\cdot N$ $0.0085(\pm 0)$ $0.0371(\pm 0.001)$ 0.0286 0.83256
Parity task.
In the following experiment, we consider a synthetic binary classification task where each input is a binary vector of length 70 and the target label is defined as the parity of 3 randomly selected input dimensions. We train a neural network using SGLD with varying values of the inverse temperature parameter $\beta$ and different sample sizes.
@@ Appendix G Mild Overparametrization Prevents Uniform Convergence
In this section, we consider fully-connected ReLU networks, where the weights are bounded, such that for each layer $j$ the absolute values of all weights are bounded by $\frac{1}{\sqrt{d_{j-1}}}$, where $d_{j-1}$ is the width of layer $j-1$. Moreover, we assume that the input $\mathbf{x}$ is such that each coordinate $x_{i}$ is bounded in $[-1,1]$. We show that $m$ training examples do not suffice for learning constant depth networks with $O(m)$ parameters. Thus, even a mild overparameterization prevents uniform convergence in our setting.
@@
We can shatter $m$ points $\{\mathbf{x}_{i}\}_{i=1}^{m}$ where $\left\|\mathbf{x}_{i}\right\|_{\infty}\leq 1$, with margin $1$, using ReLU networks of constant depth and $O(m)$ parameters, such that for each layer $j$ the absolute values of all weights are bounded by $\frac{1}{\sqrt{d_{j-1}}}$, where $d_{j-1}$ is the width of layer $j-1$.
@@
Consider input dimension $d_{0}=1$. For $1\leq i\leq m$, consider the points $x_{i}=\frac{i}{m}$, and let $\{y_{i}\}_{i=1}^{m}\subseteq\{0,1\}$. Consider the following one-hidden-layer ReLU network $N$, which satisfies $N(x_{i})=\frac{y_{i}}{m}$ for all $i$. First, the network $N$ includes a neuron with weight 00 and bias $\frac{y_{1}}{m}$, i.e., $[0\cdot x+\frac{y_{1}}{m}]_{+}$. Now, for each $i$ such that $y_{i}=0$ and $y_{i+1}=1$ we add two neurons: $[x-y_{i}]_{+}-[x-y_{i+1}]_{+}$, and for $i$ such that $y_{i}=1$ and $y_{i+1}=0$ we add $-[x-y_{i}]_{+}+[x-y_{i+1}]_{+}$. It is easy to verify that this construction has width at most $12m-$ 1 and allows us to shatter $m$ points with margin $\frac{1}{2m}$. However, the output weights of the neurons are ±$\pm 1$, and thus it does not satisfy the theorem’s requirement. Consider the network $N^{\prime}(x)=N(x)\cdot\frac{1}{\sqrt{2m-1}}$ obtained from $N$ by modifying the output weights. The network $N^{\prime}$ satisfies the theorem’s requirement on the weight magnitudes, and allows for shattering with margin $\frac{1}{2m\sqrt{2m-1}}$. We will now show how to increase this margin to $1$ using a constant number of additional layers.
Let $\tilde{N}$ be a network obtained from $N^{\prime}$ as follows. First, we add a ReLU activation to the output neuron of $N^{\prime}$. Since for every $x_{i}$ we have $N^{\prime}(x_{i})\geq 0$, it does not affect these outputs. Next, we add $L=8$ additional layers (layers $13,\ldots,3+L-$ 1) of width $\sqrt{m}$ and without bias terms, where the incoming weights to layer $3$ are all $1$ and the weights in layers $14,\ldots,3+L-$ 1 are $\frac{1}{m^{1/4}}$. Finally, we add an output neuron (layer $3+L$) with incoming weights $\frac{1}{m^{1/4}}$. The network $\tilde{N}$ satisfies the theorem’s requirements on the weight magnitudes, and it has depth $113+L=$ 11 and $O(m)$ parameters. Now, suppose that all neurons in a layer $13\leq j\leq 3+L-$ 1 have values (i.e., activations) $z\geq 0$, then the values of all neurons in layer $j+1$ are $z\cdot\frac{1}{m^{1/4}}\cdot\sqrt{m}=z\cdot m^{1/4}$. Hence, if the value of the neuron in layer $2$ is $\frac{1}{2m\sqrt{2m-1}}$, then the output of the network $\tilde{N}$ is $\frac{1}{2m\sqrt{2m-1}}\cdot(m^{1/4})^{L}=\frac{m^{L/4}}{2m\sqrt{2m-1}}=\frac{m^{2}}{2m\sqrt{2m-1}}\geq 2$ for large enough $m$. If the value of the neuron in layer $2$ is 00 then the output of $\tilde{N}$ is also 00. Hence, this construction allow for shattering $m$ points with margin at least $1$, using $O(m)$ parameters and weights that satisfy the theorem’s conditions. ∎
Appendix H Background
We supply an introduction to the theory of SDERs, then proceed to characterize the stationary distribution of a family of SDERs in a box. The background of standard (non-reflective) SDEs is similar and more common, and is therefore not included here. See for example [42] for more.
@@ H.1 SDEs with reflection
One of the main analytical tools of this work is the characterization of stationary distributions of stochastic differential equations with reflection (SDER) in bounded domains (see 44, 49, for an introduction).
The purpose of this section is to present more rigorously the setting of the paper, and supply the relevant definitions and results required to arrive at Lemma D.1. As Lemma D.1 is considered a well-known result, this section is mainly intended for completeness. Specifically, in the following we present some relevant definitions and results by Kang and Ramanan [28, 29], and specifically, ones that relate solutions to SDERs (Definition 2.4 in [29]), to solutions to sub-martingale problems (Definition 2.9 in [29]), and that characterize the stationary distributions of such solutions. For simplicity, we sometimes do not state the results in full generality.
@@
    $\eta\left(\mathbf{x}\right)=\begin{cases}\left\{\mathbf{0}\right\}&\mathbf{x}\in\Omega\,,\\ \left\{\sum_{i\in\mathcal{I}\left(\mathbf{x}\right)}\alpha_{i}{\boldsymbol{\eta}}^{i}\left(\mathbf{x}\right)\,\mid\,\alpha_{i}\geq 0\,,\;i\in\mathcal{I}\left(\mathbf{x}\right)\right\}&\mathbf{x}\in\partial\Omega\,,\end{cases}$ (36)
@@
[…]
@@
    For each $i,j\in\mathcal{I}$, $i\neq j$, and $\mathbf{x}\in\partial\Omega^{i}\cap\partial\Omega^{j}\cap\partial\Omega$,
@@
[…]
@@
Suppose that, $\Omega$ is convex and bounded, $\mathbf{b}\in\mathcal{C}^{1}\left(\overline{\Omega}\right)$ and $\mathbf{A}\in\mathcal{C}^{2}\left(\overline{\Omega}\right)$, $\left(\Omega,\eta\right)$ is piecewise $\mathcal{C}^{2}$ with continuous reflection, $\mathbf{A}$ is uniformly elliptic (see (34)), and $\mathcal{V}=\emptyset$. Then $p\in\mathcal{C}^{2}$ satisfying the conditions in Theorem H.3 defines a stationary distribution for (35).
@@
Assumptions compactness of the domain, and continuous differentiability of the drift and dispersion coefficient imply that they are Lipschitz, hence Exercise 2.5.1 and Theorem 2.5.4 of [44] imply that there exists a unique strong solution to the SDER (35). Then, piecewise $\mathcal{C}^{2}$ with continuous reflection, the uniform ellipticity assumption, Theorems 1 and 3 of [29], and Theorem H.3 imply that if there exists $p\in\mathcal{C}^{2}$ satisfying (37)-(39), then (40) is a stationary distributions of (35). ∎

### v2

@@ F.1 Experimental results
The following are results of training with SGLD (a discretized version of the CLD in (10)) on a few benchmark datasets. Notice we use the regularized version where regularization coefficient is $\lambda\cdot\beta^{-1}$ and the $\lambda$ hyperparameter is dictated by the initialization from the normal distribution $p_{0}=\mathcal{N}\left(\mathbf{0},\lambda^{-1}\mathbf{I}_{d}\right)$. We used a common initialization of $\mathcal{N}\left(\mathbf{0},\frac{1}{d_{\mathrm{in}}}\right)$, i.e. $\lambda=d_{\mathrm{in}}$.
We use several different values of $\beta$ relative to $N$ (the number of training samples). For simplicity, we focused on binary classification cases. In all datasets with more than 2 classes, we constructed a binary classification task by partitioning the original label set into $2$ disjoint sets of the same size.
The results demonstrate that learning with SGLD is possible with various values of $\beta$. In fact, in several instances, the injected noise appears to improve the generalization gap, e.g, in SVHN [53], in all the tested $\beta$ values between $0.4\cdot N$ and $2\cdot N$ the average test error remained almost the same while the training error decreased as $\beta$ increased (i.e. the generalization gap increased). Notably, we also observe that for sufficiently large levels of noise, the generalization bounds are non-vacuous.
@@
$0.01\cdot N$ $0.2279(\pm 0.0021)$ $1972(\pm 0.0243) -0.0$307 0.06124
$0.03\cdot N$ $0.1161(\pm 0.0028)$ $1074(\pm 0.0035) -0.0$087 0.10498
$1\cdot N 0.0$$618(\pm 0.001) 0.0$$62(\pm 0.0041) 0.0$002 0.19096
$15\cdot N 0.0$$497(\pm 0.0014) 0.0$$494(\pm 0.0031) -0.0$003 0.23376
$4\cdot N 0.0$$281(\pm 0.0002) 0.0$$358(\pm 0.0029) 0.0$077 0.38147
$7\cdot N 0.0$$202(\pm 0.0006) 0.0$$284(\pm 0.0024) 0.0$082 0.50456
NN $0162(\pm 0.0006) 0.0$$278(\pm 0.0023) 0.0$116 0.60302
$2\cdot N$ $0092(\pm 0.0004) 0.0$$262(\pm 0.0016) 0.0$17 0.85273
$\infty$ $0001(\pm 0) 0.0$$29(\pm 0.0004) 0.022$8 >1>1
@@
$0.01\cdot N$ $0.1215(\pm 0.0027)$ $1251(\pm 0.0087) 0.0$036 0.06833
$03\cdot N 0.0$999(±0.001)0.0999($\pm$ 0.001) $7(\pm 0.0167) 0.0088 0.11$738
$1\cdot N 0.0$$821(\pm 0.0012) 0.0$$86(\pm 0.001) 0.0$039 0.21368
$15\cdot N 0.0$$765(\pm 0.0009) 0.0$$803(\pm 0.0015) 0.0$038 0.26159
$4\cdot N 0.0$$635(\pm 0.0005) 0.0$$722(\pm 0.002) 0.0$087 0.42695
$7\cdot N 0.0$$567(\pm 0.0006) 0.0$$691(\pm 0.0019) 0.0$124 0.56473
NN $0525(\pm 0.0005) 0.0$$675(\pm 0.0013) 0.0$15 0.67495
$2\cdot N$ $043(\pm 0.0007) 0.0$$672(\pm 0.0023) 0.0$242 0.95446
$\infty$ $0248(\pm 0.001) 0.0$$675(\pm 0.0033) 0.0$427 >1>1
@@
$01\cdot N 0.0$746(±0.0012)0.0746($\pm$ 0.0012) $1033(\pm 0.0032) 0.0$287 0.05898
$03\cdot N 0.0$$441(\pm 0.0004) 0.0$$67(\pm 0.0026) 0.0$229 0.10203
$1\cdot N 0.0$$282(\pm 0.0008) 0.0$$476(\pm 0.007) 0.0$194 0.1862
$15\cdot N 0.0$$251(\pm 0.0005) 0.0$$445(\pm 0.002) 0.0$194 0.22803
$4\cdot N 0.0$$182(\pm 0.0005) 0.0$$374(\pm 0.0017) 0.0$192 0.37235
$7\cdot N 0.0$$146(\pm 0.0004) 0.0$$363(\pm 0.002) 0.0$217 0.49256
NN $0124(\pm 0.0002) 0.0$$342(\pm 0.0014) 0.0$218 0.58872
$2\cdot N$ $0085(\pm 0) 0.0$$371(\pm 0.001) 0.0$286 0.83256
@@ F.2 Training details
In this experiment, we consider a synthetic binary classification task where each input is a binary vector of length 70 and the target label is defined as the parity of 3 randomly selected input dimensions. We train a neural network using SGLD with varying values of the inverse temperature parameter $\beta$ and different sample sizes.
@@
F.3 Comparison with the bound of Mou et al. [49]
The bound proposed by Mou et al. [49] has demonstrated non-vacuous results. To further assess the effectiveness of our bound and evaluate its relative tightness, we conducted a series of numerical experiments on the MNIST binary classification task (see Tables 5-8).
It is worth emphasizing that our bound offers a distinct advantage: it can be evaluated directly at initialization, whereas the bound of Mou et al. [49] depends on gradients and therefore cannot be computed before training. When testing their bound we used the continuous version, i.e.
$\mathbb{E}_{p_{T}}\!\left[E_{D}\!\left({\boldsymbol{\theta}}\right)\!-E_{S}\!\left({\boldsymbol{\theta}}\right)\right]\leq s\left(\frac{\beta}{2n}\int_{0}^{T}e^{\frac{\lambda}{2}(T-t)}\mathbb{E}_{p_{t}}\!\left[\|\nabla L_{S}({\boldsymbol{\theta}})\|^{2}\right]dt+\frac{\log(1/\delta)+\log\log M}{n}\right)^{0.5}\,.$
For simplicity, we omitted the term involving $M$ (which makes the bound more favorable). In addition, we set $s=0.5$ since the zero–one loss (denoted here by $f(w)$, unlike [49]) is bounded within the interval $[0,1]$. We observed that the relative tightness of the two bounds varies across different values of $\beta$ and at different points in time. Consequently, in some instances, the bound of Mou et al. [49] is tighter, while in others our bound performs better, and we could not draw any further conclusions.
Table 5: 20 training epochs
$\beta$ Train Error Test Error Generalization Gap Mou et al. [49] Our bound
$0.03N=1800$ $0.1224$ $0.137$ $0.0146$ $0.0539$ $0.1144$
$0.15N=9000$ $0.0515$ $0.0747$ $0.0232$ $0.1279$ $0.2548$
$0.4N=24000$ $0.0335$ $0.058$ $0.0245$ $0.2845$ $0.4157$
$0.7N=42000$ $0.0278$ $0.0498$ $0.0220$ $0.4930$ $0.5499$
$N=60000$ $0.0249$ $0.0428$ $0.0179$ $0.7032$ $0.6572$
$2N=120000$ $0.0209$ $0.0356$ $0.0147$ $1.4044$ $0.9294$
Table 6: 50 training epochs
$\beta$ Train Error Test Error Generalization Gap Mou et al. [49] Our bound
$0.03N=1800$ $0.1156$ $0.1697$ $0.0541$ $0.0637$ $0.1144$
$0.15N=9000$ $0.0491$ $0.0615$ $0.0124$ $0.1324$ $0.2548$
$0.4N=24000$ $0.0295$ $0.0348$ $0.0053$ $0.2992$ $0.4157$
$0.7N=42000$ $0.0217$ $0.0283$ $0.0066$ $0.4903$ $0.5499$
$N=60000$ $0.0173$ $0.0277$ $0.0104$ $0.6827$ $0.6572$
$2N=120000$ $0.0108$ $0.0265$ $0.0157$ $1.3153$ $0.9294$
Table 7: 250 training epochs
$\beta$ Train Error Test Error Generalization Gap Mou et al. [49] Our bound
$0.03N=1800$ $0.122$ $0.1049$ 0.0171-0.0171 $0.1273$ $0.1144$
$0.15N=9000$ $0.0502$ $0.0476$ 0.0026-0.0026 $0.1503$ $0.2548$
$0.4N=24000$ $0.0284$ $0.0296$ $0.0011$ $0.2853$ $0.4157$
$0.7N=42000$ $0.0178$ $0.0247$ $0.0069$ $0.4595$ $0.5499$
$N=60000$ $0.0127$ $0.0240$ $0.0113$ $0.6478$ $0.6572$
$2N=120000$ $0.0050$ $0.0234$ $0.0184$ $1.2158$ $0.9294$
Table 8: 400 training epochs
$\beta$ Train Error Test Error Generalization Gap Mou et al. [49] Our bound
$0.03N=1800$ $0.1224$ $0.1105$ 0.0119-0.0119 $0.1900$ $0.1144$
$0.15N=9000$ $0.0499$ $0.0556$ $0.0057$ $0.1774$ $0.2548$
$0.4N=24000$ $0.0261$ $0.0357$ $0.0096$ $0.3005$ $0.4157$
$0.7N=42000$ $0.0161$ $0.0271$ $0.0110$ $0.4548$ $0.5499$
$N=60000$ $0.0112$ $0.0255$ $0.0143$ $0.6247$ $0.6572$
$2N=120000$ $0.0038$ $0.0249$ $0.0211$ $1.1455$ $0.9294$
@@ Appendix G Mild Overparametrization Prevents Uniform Convergence
In this section, we consider fully-connected ReLU networks, where the weights are bounded, such that for each layer $j$ the absolute values of all weights are bounded by $\frac{1}{\sqrt{d_{j-1}}}$, where $d_{j-1}$ is the width of layer j1j-1. Moreover, we assume that the input $\mathbf{x}$ is such that each coordinate $x_{i}$ is bounded in [1,1][-1,1]. We show that $m$ training examples do not suffice for learning constant depth networks with $O(m)$ parameters. Thus, even a mild overparameterization prevents uniform convergence in our setting.
@@
We can shatter $m$ points $\{\mathbf{x}_{i}\}_{i=1}^{m}$ where $\left\|\mathbf{x}_{i}\right\|_{\infty}\leq 1$, with margin $1$, using ReLU networks of constant depth and $O(m)$ parameters, such that for each layer $j$ the absolute values of all weights are bounded by $\frac{1}{\sqrt{d_{j-1}}}$, where $d_{j-1}$ is the width of layer j1j-1.
@@
Consider input dimension $d_{0}=1$. For $1\leq i\leq m$, consider the points $x_{i}=\frac{i}{m}$, and let $\{y_{i}\}_{i=1}^{m}\subseteq\{0,1\}$. Consider the following one-hidden-layer ReLU network $N$, which satisfies $N(x_{i})=\frac{y_{i}}{m}$ for all $i$. First, the network $N$ includes a neuron with weight 0 and bias $\frac{y_{1}}{m}$, i.e., $0\cdot x+\frac{y_{1}}{m}]_{+}$. Now, for each $i$ such that $y_{i}=0$ and $y_{i+1}=1$ we add two neurons: $x-y_{i}]_{+}-[x-y_{i+1}]_{+}$, and for $i$ such that $y_{i}=1$ and $y_{i+1}=0$ we add $x-y_{i}]_{+}+[x-y_{i+1}]_{+}$. It is easy to verify that this construction has width at most 2m12m-1 and allows us to shatter $m$ points with margin $\frac{1}{2m}$. However, the output weights of the neurons are $\pm 1$, and thus it does not satisfy the theorem’s requirement. Consider the network $N^{\prime}(x)=N(x)\cdot\frac{1}{\sqrt{2m-1}}$ obtained from $N$ by modifying the output weights. The network $N^{\prime}$ satisfies the theorem’s requirement on the weight magnitudes, and allows for shattering with margin $\frac{1}{2m\sqrt{2m-1}}$. We will now show how to increase this margin to $1$ using a constant number of additional layers.
Let $\tilde{N}$ be a network obtained from $N^{\prime}$ as follows. First, we add a ReLU activation to the output neuron of $N^{\prime}$. Since for every $x_{i}$ we have $N^{\prime}(x_{i})\geq 0$, it does not affect these outputs. Next, we add $L=8$ additional layers (layers $3,\ldots,3+L-1$) of width $\sqrt{m}$ and without bias terms, where the incoming weights to layer $3$ are all $1$ and the weights in layers $4,\ldots,3+L-1$ are $\frac{1}{m^{1/4}}$. Finally, we add an output neuron (layer $3+L$) with incoming weights $\frac{1}{m^{1/4}}$. The network $\tilde{N}$ satisfies the theorem’s requirements on the weight magnitudes, and it has depth $3+L=11$ and $O(m)$ parameters. Now, suppose that all neurons in a layer $3\leq j\leq 3+L-1$ have values (i.e., activations) $z\geq 0$, then the values of all neurons in layer $j+1$ are $z\cdot\frac{1}{m^{1/4}}\cdot\sqrt{m}=z\cdot m^{1/4}$. Hence, if the value of the neuron in layer $2$ is $\frac{1}{2m\sqrt{2m-1}}$, then the output of the network $\tilde{N}$ is $\frac{1}{2m\sqrt{2m-1}}\cdot(m^{1/4})^{L}=\frac{m^{L/4}}{2m\sqrt{2m-1}}=\frac{m^{2}}{2m\sqrt{2m-1}}\geq 2$ for large enough $m$. If the value of the neuron in layer $2$ is 0 then the output of $\tilde{N}$ is also 0. Hence, this construction allow for shattering $m$ points with margin at least $1$, using $O(m)$ parameters and weights that satisfy the theorem’s conditions. ∎
Appendix H Background on Stochastic Differential Equations with Reflection
We supply an introduction to the theory of stochastic differential equations with reflection (SDERs), then proceed to characterize the stationary distribution of a family of SDERs in a box. The background of standard (non-reflective) SDEs is similar and more common, and is therefore not included here. See for example [54] for more.
@@ H.1 SDEs with reflection
One of the main analytical tools of this work is the characterization of stationary distributions of SDER in bounded domains (see 57, 63, for an introduction).
The purpose of this section is to present more rigorously the setting of the paper, and supply the relevant definitions and results required to arrive at Lemma˜D.1. As Lemma˜D.1 is considered a well-known result, this section is mainly intended for completeness. Specifically, in the following we present some relevant definitions and results by Kang and Ramanan [31, 32], and specifically, ones that relate solutions to SDERs (Definition 2.4 in [32]), to solutions to sub-martingale problems (Definition 2.9 in [32]), and that characterize the stationary distributions of such solutions. For simplicity, we sometimes do not state the results in full generality.
@@
[…]
@@
Suppose that the pair $\left(\Omega,\eta\right)$ is piecewise $\mathcal{C}^{2}$ with continuous reflection, for all $i\in\mathcal{I}$ and $\mathbf{x}\in\partial\Omega^{i}$, $\left\langle\mathbf{n}^{i}\left(\mathbf{x}\right),{\boldsymbol{\eta}}^{i}\left(\mathbf{x}\right)\right\rangle=1$, $\mathcal{V}=\emptyset$, $\mathbf{b}\left(\cdot\right)\in\mathcal{C}^{1}\left(\overline{\Omega}\right)$ and $\mathbf{A}\in\mathcal{C}^{2}\left(\overline{\Omega}\right)$ (elementwise), and the submartingale problem associated with $\left(\Omega,\eta\right)$ and $\mathcal{V}$ is well posed. Furthermore, suppose there exists a nonnegative function $p\in\mathcal{C}^{2}\left(\overline{\Omega}\right)$ with $Z_{p}=\int_{\overline{\Omega}}p\left(\mathbf{x}\right)d\mathbf{x}<\infty$ that solves the PDE defined by the following three relations:
@@
    For each $\in\mathcal{I}, iji\neq j$, and $\mathbf{x}\in\partial\Omega^{i}\cap\partial\Omega^{j}\cap\partial\Omega$,
@@
$p_{\infty}\left(A\right)\triangleq\frac{1}{Z_{p}}\intop_{A}p\left(\mathbf{x}\right)d\mathbf{x}\,,\quad A\in\mathcal{B}\left(\overline{\Omega}\right)\,$, (44)
@@
Suppose that, $\Omega$ is convex and bounded, $\mathbf{b}\in\mathcal{C}^{1}\left(\overline{\Omega}\right)$ and $\mathbf{A}\in\mathcal{C}^{2}\left(\overline{\Omega}\right)$, $\left(\Omega,\eta\right)$ is piecewise $\mathcal{C}^{2}$ with continuous reflection, $\mathbf{A}$ is uniformly elliptic (see (38)), and $\mathcal{V}=\emptyset$. Then $p\in\mathcal{C}^{2}$ satisfying the conditions in Theorem˜H.3 defines a stationary distribution for (39).
@@
Assumptions compactness of the domain, and continuous differentiability of the drift and dispersion coefficient imply that they are Lipschitz, hence Exercise 2.5.1 and Theorem 2.5.4 of [57] imply that there exists a unique strong solution to the SDER (39). Then, piecewise $\mathcal{C}^{2}$ with continuous reflection, the uniform ellipticity assumption, Theorems 1 and 3 of [32], and Theorem˜H.3 imply that if there exists $p\in\mathcal{C}^{2}$ satisfying (41)-(43), then (44) is a stationary distributions of (39). ∎